# EAG3R: Event-Augmented 3D Geometry Estimation for Dynamic and Extreme-Lighting Scenes

### v1

@@ 1 Introduction
Estimating geometry from videos or images is a fundamental problem in 3D vision, with broad applications in camera pose estimation, novel view synthesis, geometry reconstruction, and 3D perception. These capabilities are crucial in downstream scenarios such as autonomous driving, SLAM, virtual environments, and robotic navigation. Recent methods like DUSt3R wang2024dust3r have shown that regressing dense pointmaps from image pairs using transformer-based foundation models enables accurate and efficient pose-free 3D reconstruction. This paradigm has sparked a growing trend toward addressing various challenging scenarios, such as longer image sequences wang20243d; wang2025continuous; wang2025vggt, dynamic scenes zhang2024monst3r; chen2025easi3r; jin2024stereo4d; sucar2025dynamic, and integration with techniques like Gaussian Splatting fan2024instantsplat; smart2024splatt3r; gao2025easysplat.
@@
Event cameras, on the other hand, provide asynchronous measurements of pixel-level brightness changes with high temporal resolution and dynamic range. They have demonstrated strong resilience in challenging conditions such as fast motion and extreme illumination Gallego2022Event; Kim2016RealTime; Rebecq2018EMVS. Prior work has leveraged event streams in 3D tasks such as depth estimation Baudron2020E3D; Zuo2022DEVO; Nam2022StereoDepthEvents, surface reconstruction Carneiro2013EventBased3D; Chen2023DenseVoxel, and neural rendering Rudnev2023EventNeRF; huang2024inceventgs, but their integration into modern learning-based geometry pipelines remains limited.
In this paper, we propose EAG3R, an event-augemented MonST3R framework to enhance pointmap-based 3D geometry estimation under dynamic and extremely low-light conditions. Built upon the MonST3R zhang2024monst3r backbone, EAG3R introduces two key innovations: (1) a lightweight event adapter with Signal-to-Noise Ratio (SNR)-aware fusion mechanism that adaptively integrates event and image features based on local reliability, and (2) an event-based photometric consistency loss that enforces alignment between predicted motion-induced brightness changes and event-observed brightness changes during global optimization. These components enable EAG3R to remain robust in scenarios where conventional RGB-only pipelines fail.
We evaluate EAG3R on the MVSEC dataset zhu2018multivehicle, conducting extensive comparisons on depth estimation, camera pose tracking, and dynamic reconstruction in extreme low-light conditions. Results show that EAG3R significantly outperforms existing baselines, including DUSt3R wang2024dust3r, MonST3R zhang2024monst3r, and Easi3R chen2025easi3r variants, even in a zero-shot nighttime setting.
@@ 2 Related Work
Traditional Structure-from-Motion (SfM) agarwal2011building; pollefeys1999self; pollefeys2004visual; schonberger2016structure; snavely2006photo; snavely2008modeling and Simultaneous Localization and Mapping (SLAM) davison2007monoslam; engel2014lsd; mur2015orb; newcombe2011dtam methods estimate 3D structure and camera motion by establishing 2D correspondences bay2008speeded; de2018superpoint; lowe2004distinctive; mur2015orb; sarlin2020superglue or minimizing photometric errors engel2017direct; engel2014lsd, followed by bundle adjustment (BA) agarwal2010bundle; brachmann2024scene; wang2024vggsfm. While effective with dense views, these often struggle with sparse or ill-conditioned data. Recent learning-based approaches aim to improve robustness and efficiency: DUSt3R wang2024dust3r, directly regress dense point maps from image pairs using Transformer architectures dosovitskiy2020image trained on large-scale 3D datasets. However, DUSt3R and its variants liu2024slam3r; murai2024mas13r; tang2024mv; wang20243d; wang2025vggt are primarily designed for static scenes and their performance degrades with dynamic content.
@@
Reconstructing dynamic scenes without known poses is a core challenge in SLAM. Classical methods rely on joint pose estimation and dynamic region filtering via semantics yu2018ds or optical flow zhao2022particlesfm, but depend heavily on accurate segmentation and tracking. Some methods estimate temporally consistent depth using geometric constraints luo2020consistent or generative priors hu2024depthcrafter; shao2024learning, yet suffer from fragmented reconstructions due to missing camera trajectories. Recent works jointly optimize depth and pose by refining pretrained depth models ranftl2020towards using flow wang2024sea and masks he2017mask, e.g., in CasualSAM zhang2022structure, Robust-CVD kopf2021robust, and MegaSAM li2024megasam, the latter integrating DROID-SLAM teed2021droid and diverse priors piccinelli2024unidepth; yang2024depth. Recent approaches employ direct pointmap regression, including MonST3R zhang2024monst3r, DAS3R xu2024das3r, CUT3R wang2025continuous, and Easi3R chen2025easi3r, which leverage optical flow, segmentation, or attention for motion disentanglement.
@@
Low-light image enhancement (LIE) aims to improve image quality under poor illumination. Traditional methods like histogram equalization Arici2009 and Retinex-based algorithms Guo2016 have limited adaptability, while deep learning approaches Wei2018; cai2023retinexformer; Wang2019Learning achieve better results but still struggle in extreme darkness. Event cameras, with high dynamic range and temporal resolution, enable structural information preservation in very low light Zheng2023Survey; Zhang2020Events, inspiring event-guided LIE methods Jiang2023Event; Liu2023SyntheticEvent. However, robust fusion of frame and event data under noise remains challenging. EvLight Liang2024EvLight addresses this with adaptive event-image feature fusion.
@@
Event cameras have enabled progress in 3D reconstruction under challenging lighting and fast motion Gallego2022Event. Early work used stereo setups for disparity and multi-view stereo Carneiro2013EventBased3D; Piatkowska2013AsynchronousStereo; Zhu2018Realtime, followed by monocular methods based on geometric priors like camera trajectories Kim2016RealTime; Rebecq2018EMVS. Recent approaches apply deep learning to stereo Nam2022StereoDepthEvents and monocular Baudron2020E3D; Chen2023DenseVoxel settings, producing dense outputs such as meshes or voxels. Multimodal fusion with structured light Leroux2018EventBasedStructuredLight or RGB-D sensors Zuo2022DEVO further improves robustness. Latest advances adapt NeRF Klenk2023ENeRF; Rudnev2023EventNeRF and Gaussian Splatting han2024event; huang2024inceventgs to event data, enabling high-fidelity scene reconstruction and novel view synthesis.
@@ 3.1 Preliminary
However, the reliance of these image-based methods on clear visual information causes them to struggle in low-light settings: RGB images $I^{t}$ lose crucial detail, and MonST3R’s flow estimation with RAFT  teed2020raft can become unstable. Our EAG3R addresses these limitations by integrating asynchronous event data $E^{t}$ into both the feature extraction and global optimization stages, aiming for robust 3D reconstruction performance even in such challenging lighting conditions.
@@ 3.2 Event-data Integration and Feature Fusion
To enhance image visibility under low-light conditions and provide a spatial reliability prior, we introduce a lightweight Retinex-inspired cai2023retinexformer enhancement module. Given an input image $I^{t}$, we estimate an illumination map $L_{\text{illum}}^{t}$ using a shallow network $\mathcal{N}_{\text{retinex}}$ with inputs $I^{t}$ and its channel-wise maximum projection $I_{\text{prior}}^{t}=\max_{c}(I^{t})$, and compute the enhanced image via element-wise multiplication:
@@
To effectively harness asynchronous event streams $\mathcal{E}^{t}$ for dense geometric prediction, we introduce a lightweight event adapter by employing a Swin Transformer backbone initialized with weights from a self-supervised, context-based pre-training regimen on event data yang2023event. The input events from $\mathcal{E}^{t}$ are voxelized into a spatiotemporal grid and processed by the pre-trained Swin Transformer encoder, yielding hierarchical event features $\{F_{\text{evt},l}^{t}\}_{l=1}^{4}$. Corresponding hierarchical image features $\{F_{\text{img},l}^{t}\}_{l=1}^{4}$ are extracted at every 6 layers from intermediate representations within the pre-trained image encoder. At each hierarchical stage $l$, the event features $F_{\text{evt},l}^{t}$ are spatially aligned and dimension-matched with their respective image counterparts $F_{\text{img},l}^{t}$.
We then apply cross-attentionxu2022snr, using event features as queries and image features as keys and values:
@@ 3.3.1 Event-Based Photometric Consistency Loss
where $\Delta\tau$ denotes the integration interval and $C$ is the contrast sensitivity threshold intrinsic to the event sensor. This expression follows the generative model introduced in Gehrig2018eccv.
@@ 4 Experiments
We evaluate our method in a variety of tasks, including depth estimation (Section 4.2), camera pose estimation (Section 4.3) and 4D reconstruction (Section 4.4). We perform ablation studies in Section 4.5. We compare EAG3R with state-of-the-art pose free learning-based reconstruction method, including DUSt3R wang2024dust3r, MonST3R zhang2024monst3r, and Easi3R chen2025easi3r.
@@ 4.1 Experiment Details
For dataset selection, we initially attempted to fine-tune the MonST3R baseline using events generated via Video-to-Events (V2E)Gehrig_2020_CVPR from MonST3R’s fine-tuning datasets. However, the noise in V2E-generated events led to gradient explosion during training, prompting us to switch to datasets with real event captures and ground truth (GT) depth. Given the scarcity of such data, we selected the Multi Vehicle Stereo Event Camera (MVSEC) dataset zhu2018multivehicle. It provides synchronized stereo events and reliable LiDAR-derived depth GT. To ensure a fair zero-shot evaluation of low-light performance, all models were trained exclusively on MVSEC’s outdoor_day2 sequence (normal daylight) and tested on the challenging outdoor_night1-3 sequences (extreme low-light).
@@ 4.2 Monocular Depth Estimation
DUSt3Rwang2024dust3r 0.407 0.393 0.534 0.415 0.384 0.495 0.463 0.335 0.534
MonST3Rzhang2024monst3r 0.370 0.373 0.497 0.309 0.469 0.409 0.317 0.453 0.418
@@ 4.3 Camera Pose Estimation
DUSt3Rwang2024dust3r 1.474 0.914 2.995 3.921 2.207 10.761 4.109 2.401 11.309
MonST3Rzhang2024monst3r 0.559 0.317 0.369 0.626 0.341 1.460 0.733 0.427 1.371
@@ A.2 Video Depth Estimation Results on MVSEC
DUSt3Rwang2024dust3r 0.432 0.374 0.547 0.410 0.397 0.493 0.510 0.322 0.554
MonST3Rzhang2024monst3r 0.380 0.388 0.486 0.299 0.494 0.388 0.296 0.499 0.392
@@ A.5.4 Feature Strategy for Global Optimization
SuperPoint de2018superpoint 0.685 0.260 0.153 High
@@ A.9 Limitations
Our approach assumes access to temporally aligned, high-quality event streams. Although the SNR-aware fusion mechanism mitigates some effects of noise, the performance of EAG3R can still degrade when event data are excessively sparse, noisy, or misaligned. In particular, we attempted to train our model using synthetic events generated by V2E Gehrig2018eccv, but observed that the low fidelity of these generated events caused optimization instability, including gradient explosion and failure to converge.

### v2

@@ 1 Introduction
Estimating geometry from videos or images is a fundamental problem in 3D vision, with broad applications in camera pose estimation, novel view synthesis, geometry reconstruction, and 3D perception. These capabilities are crucial in downstream scenarios such as autonomous driving, SLAM, virtual environments, and robotic navigation. Recent methods like DUSt3R Wang et al. (2024b) have shown that regressing dense pointmaps from image pairs using transformer-based foundation models enables accurate and efficient pose-free 3D reconstruction. This paradigm has sparked a growing trend toward addressing various challenging scenarios, such as longer image sequences Wang and Agapito (2024); Wang et al. (2025b, a), dynamic scenes Zhang et al. (2024); Chen et al. (2025); Jin et al. (2024); Sucar et al. (2025), and integration with techniques like Gaussian Splatting Fan et al. (2024); Smart et al. (2024); Gao et al. (2025).
@@
Event cameras, on the other hand, provide asynchronous measurements of pixel-level brightness changes with high temporal resolution and dynamic range. They have demonstrated strong resilience in challenging conditions such as fast motion and extreme illumination Gallego et al. (2022); Kim et al. (2016); Rebecq et al. (2018). Prior work has leveraged event streams in 3D tasks such as depth estimation Baudron et al. (2020); Zuo et al. (2022); Nam et al. (2022), surface reconstruction Carneiro et al. (2013); Chen et al. (2023), and neural rendering Rudnev et al. (2023); Huang et al. (2025), but their integration into modern learning-based geometry pipelines remains limited.
In this paper, we propose EAG3R, an event-augemented MonST3R framework to enhance pointmap-based 3D geometry estimation under dynamic and extremely low-light conditions. Built upon the MonST3R Zhang et al. (2024) backbone, EAG3R introduces two key innovations: (1) a lightweight event adapter with Signal-to-Noise Ratio (SNR)-aware fusion mechanism that adaptively integrates event and image features based on local reliability, and (2) an event-based photometric consistency loss that enforces alignment between predicted motion-induced brightness changes and event-observed brightness changes during global optimization. These components enable EAG3R to remain robust in scenarios where conventional RGB-only pipelines fail.
We evaluate EAG3R on the MVSEC dataset Zhu et al. (2018b), conducting extensive comparisons on depth estimation, camera pose tracking, and dynamic reconstruction in extreme low-light conditions. Results show that EAG3R significantly outperforms existing baselines, including DUSt3R Wang et al. (2024b), MonST3R Zhang et al. (2024), and Easi3R Chen et al. (2025) variants, even in a zero-shot nighttime setting.
@@ 2 Related Work
Traditional Structure-from-Motion (SfM) Agarwal et al. (2011); Pollefeys et al. (1999, 2004); Schönberger and Frahm (2016); Snavely et al. (2006, 2008) and Simultaneous Localization and Mapping (SLAM) Davison et al. (2007); Engel et al. (2014); Mur-Artal et al. (2015); Newcombe et al. (2011) methods estimate 3D structure and camera motion by establishing 2D correspondences Bay et al. (2008); DeTone et al. (2018); Lowe (2004); Mur-Artal et al. (2015); Sarlin et al. (2020) or minimizing photometric errors Engel et al. (2017, 2014), followed by bundle adjustment (BA) Agarwal et al. (2010); Brachmann et al. (2024); Wang et al. (2024a). While effective with dense views, these often struggle with sparse or ill-conditioned data. Recent learning-based approaches aim to improve robustness and efficiency: DUSt3R Wang et al. (2024b), directly regress dense point maps from image pairs using Transformer architectures Dosovitskiy et al. (2020) trained on large-scale 3D datasets. However, DUSt3R and its variants Liu et al. (2025); Murai et al. (2025); Tang et al. (2025); Wang and Agapito (2024); Wang et al. (2025a) are primarily designed for static scenes and their performance degrades with dynamic content.
@@
Reconstructing dynamic scenes without known poses is a core challenge in SLAM. Classical methods rely on joint pose estimation and dynamic region filtering via semantics Yu et al. (2018) or optical flow Zhao et al. (2022), but depend heavily on accurate segmentation and tracking. Some methods estimate temporally consistent depth using geometric constraints Luo et al. (2020) or generative priors Hu et al. (2025); Shao et al. (2025), yet suffer from fragmented reconstructions due to missing camera trajectories. Recent works jointly optimize depth and pose by refining pretrained depth models Ranftl et al. (2020) using flow Wang et al. (2024c) and masks He et al. (2017), e.g., in CasualSAM Zhang et al. (2022), Robust-CVD Kopf et al. (2021), and MegaSAM Li et al. (2025), the latter integrating DROID-SLAM Teed and Deng (2021) and diverse priors Piccinelli et al. (2024); Yang et al. (2024). Recent approaches employ direct pointmap regression, including MonST3R Zhang et al. (2024), DAS3R Xu et al. (2024), CUT3R Wang et al. (2025b), and Easi3R Chen et al. (2025), which leverage optical flow, segmentation, or attention for motion disentanglement.
@@
Low-light image enhancement (LIE) aims to improve image quality under poor illumination. Traditional methods like histogram equalization Arici et al. (2009) and Retinex-based algorithms Guo et al. (2016) have limited adaptability, while deep learning approaches Wei et al. (2018); Cai et al. (2023); Wang et al. (2019) achieve better results but still struggle in extreme darkness. Event cameras, with high dynamic range and temporal resolution, enable structural information preservation in very low light Zheng et al. (2023); Zhang et al. (2020), inspiring event-guided LIE methods Jiang et al. (2023); Liu et al. (2023). However, robust fusion of frame and event data under noise remains challenging. EvLight Liang et al. (2024) addresses this with adaptive event-image feature fusion.
@@
Event cameras have enabled progress in 3D reconstruction under challenging lighting and fast motion Gallego et al. (2022). Early work used stereo setups for disparity and multi-view stereo Carneiro et al. (2013); Piatkowska et al. (2013); Zhu et al. (2018a), followed by monocular methods based on geometric priors like camera trajectories Kim et al. (2016); Rebecq et al. (2018). Recent approaches apply deep learning to stereo Nam et al. (2022) and monocular Baudron et al. (2020); Chen et al. (2023) settings, producing dense outputs such as meshes or voxels. Multimodal fusion with structured light Leroux et al. (2018) or RGB-D sensors Zuo et al. (2022) further improves robustness. Latest advances adapt NeRF Klenk et al. (2023); Rudnev et al. (2023) and Gaussian Splatting Han et al. (2024); Huang et al. (2025) to event data, enabling high-fidelity scene reconstruction and novel view synthesis.
@@ 3.1 Preliminary
However, the reliance of these image-based methods on clear visual information causes them to struggle in low-light settings: RGB images $I^{t}$ lose crucial detail, and MonST3R’s flow estimation with RAFT  Teed and Deng (2020) can become unstable. Our EAG3R addresses these limitations by integrating asynchronous event data $E^{t}$ into both the feature extraction and global optimization stages, aiming for robust 3D reconstruction performance even in such challenging lighting conditions.
@@ 3.2 Event-data Integration and Feature Fusion
To enhance image visibility under low-light conditions and provide a spatial reliability prior, we introduce a lightweight Retinex-inspired Cai et al. (2023) enhancement module. Given an input image $I^{t}$, we estimate an illumination map $L_{\text{illum}}^{t}$ using a shallow network $\mathcal{N}_{\text{retinex}}$ with inputs $I^{t}$ and its channel-wise maximum projection $I_{\text{prior}}^{t}=\max_{c}(I^{t})$, and compute the enhanced image via element-wise multiplication:
@@
To effectively harness asynchronous event streams $\mathcal{E}^{t}$ for dense geometric prediction, we introduce a lightweight event adapter by employing a Swin Transformer backbone initialized with weights from a self-supervised, context-based pre-training regimen on event data Yang et al. (2023). The input events from $\mathcal{E}^{t}$ are voxelized into a spatiotemporal grid and processed by the pre-trained Swin Transformer encoder, yielding hierarchical event features $\{F_{\text{evt},l}^{t}\}_{l=1}^{4}$. Corresponding hierarchical image features $\{F_{\text{img},l}^{t}\}_{l=1}^{4}$ are extracted at every 6 layers from intermediate representations within the pre-trained image encoder. At each hierarchical stage $l$, the event features $F_{\text{evt},l}^{t}$ are spatially aligned and dimension-matched with their respective image counterparts $F_{\text{img},l}^{t}$.
We then apply cross-attentionXu et al. (2022), using event features as queries and image features as keys and values:
@@ 3.3.1 Event-Based Photometric Consistency Loss
where $\Delta\tau$ denotes the integration interval and $C$ is the contrast sensitivity threshold intrinsic to the event sensor. This expression follows the generative model introduced in Gehrig et al. (2018).
@@ 4 Experiments
We evaluate our method in a variety of tasks, including depth estimation (Section 4.2), camera pose estimation (Section 4.3) and 4D reconstruction (Section 4.4). We perform ablation studies in Section 4.5. We compare EAG3R with state-of-the-art pose free learning-based reconstruction method, including DUSt3R Wang et al. (2024b), MonST3R Zhang et al. (2024), and Easi3R Chen et al. (2025).
@@ 4.1 Experiment Details
For dataset selection, we initially attempted to fine-tune the MonST3R baseline using events generated via Video-to-Events (V2E)Gehrig et al. (2020) from MonST3R’s fine-tuning datasets. However, the noise in V2E-generated events led to gradient explosion during training, prompting us to switch to datasets with real event captures and ground truth (GT) depth. Given the scarcity of such data, we selected the Multi Vehicle Stereo Event Camera (MVSEC) dataset Zhu et al. (2018b). It provides synchronized stereo events and reliable LiDAR-derived depth GT. To ensure a fair zero-shot evaluation of low-light performance, all models were trained exclusively on MVSEC’s outdoor_day2 sequence (normal daylight) and tested on the challenging outdoor_night1-3 sequences (extreme low-light).
@@ 4.2 Monocular Depth Estimation
DUSt3RWang et al. (2024b) 0.407 0.393 0.534 0.415 0.384 0.495 0.463 0.335 0.534
MonST3RZhang et al. (2024) 0.370 0.373 0.497 0.309 0.469 0.409 0.317 0.453 0.418
@@ 4.3 Camera Pose Estimation
DUSt3RWang et al. (2024b) 1.474 0.914 2.995 3.921 2.207 10.761 4.109 2.401 11.309
MonST3RZhang et al. (2024) 0.559 0.317 0.369 0.626 0.341 1.460 0.733 0.427 1.371
@@ A.2 Video Depth Estimation Results on MVSEC
DUSt3RWang et al. (2024b) 0.432 0.374 0.547 0.410 0.397 0.493 0.510 0.322 0.554
MonST3RZhang et al. (2024) 0.380 0.388 0.486 0.299 0.494 0.388 0.296 0.499 0.392
@@ A.5.4 Feature Strategy for Global Optimization
SuperPoint DeTone et al. (2018) 0.685 0.260 0.153 High
@@ A.9 Limitations
Our approach assumes access to temporally aligned, high-quality event streams. Although the SNR-aware fusion mechanism mitigates some effects of noise, the performance of EAG3R can still degrade when event data are excessively sparse, noisy, or misaligned. In particular, we attempted to train our model using synthetic events generated by V2E Gehrig et al. (2018), but observed that the low fidelity of these generated events caused optimization instability, including gradient explosion and failure to converge.